OpenStudy (anonymous):

If the midpoint between (8, y) and (-11, 6) is (-1.5, 5), find the value of y

OpenStudy (campbell_st):

well using the information 5 = ( 6 + y)/2 solve for y

OpenStudy (anonymous):

how would i do that

OpenStudy (anonymous):

to answer 5=(6+y)/2 you can use this... http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

its pretty self explanitory

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

@campbell_st how did you get that formula

OpenStudy (anonymous):

@Walkerman15

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

5 = ( 6 + y)/2 how would i find the formula he gave me

OpenStudy (campbell_st):

the midpoint... is simply the average of the x values and average of the y values. There is a formula... but it's easier to remember average x and average y so add the 2 values and divide the answer by 2 hope it makes sense